### v1

@@ 5.1.1 Answer to RQ1
Using passive voice in natural language requirements specifications has a slight negative effect on the domain modeling activity regarding missing associations. This finding aligns with the conclusions drawn by the original study by Femmer et al. femmer2014impact . However, the Bayesian data analysis emphasizes that both context factors, and especially the number of missing entities, have a significant impact on the number of missing associations as well. Overall, these results support the claim that passive voice can have a negative impact in specific cases but is overall not a significant factor in subsequent activities depending on the requirement femmer2014impact ; krisch2015myth .
@@ 5.2 Implications for Requirements Quality Practice
Finally, the potential influence of context factors on the impact of quality defects on affected activities may incentivize organizations to invest in developing these factors. For example, improving domain knowledge and providing formal modeling training may compensate quality defects.
@@ 5.3 Implications for Requirements Quality Research
Employing Bayesian data analysis to investigate the impact of requirements quality defects provides sophisticated and sensitive insights necessary to propel requirements quality research frattini2023requirements . The result of the analysis models both the direction and strength of an impacting factor while retaining information about its certainty. These insights go beyond the point-wise comparison and binary result of frequentist analyses furia2019bayesian . The frequentist analysis fails to compare the impact of quality defects on the response variables, as even the calculated effect sizes are similar ($0.930.79<|ES|<$ 0.93). The Bayesian data analysis, on the other hand, clearly shows that some effects are much stronger (e.g., $H_{0}^{AP\rightarrow A^{\times}}$) than others (e.g., $H_{0}^{PV\rightarrow A^{-}}$).
Including context factors in the prediction allows comparing the impact of requirements quality with the impact of human and process factors, revealing which causes changes in the response variable. Conducting these studies justifies tool development. Additionally, it increases the likelihood of solving problems with practical relevance franch2020practitioners . However, involving context factors in the evaluation requires accurate metrics. In this study, we followed an approach common in SE research by measuring context factors via self-assessment on a discrete scale. Better metrics of context factors will allow modeling their influence more accurately.
@@
Using a controlled experiment benefits the investigation of the quality factor femmer2014impact . The DAG shown in Figure 9 visualizes this control, as other factors do not influence the treatment in question. This eliminates spurious associations that could confuse the results mcelreath2020statistical . On the other hand, the cost of conducting a controlled experiment—especially with participants from industry—cannot be neglected sjoberg2003challenges . Luckily, statistical causal inference via Bayesian data analysis works equally well with observational data, as shown by Furia et al. furia2023towards .
Worth mentioning is also the synergy between Bayesian data analysis and experiments with a crossover design. The main threats to validity in crossover designs include the impact of confounding variables like the learning and crossover effect kitchenham2003case . The strategies proposed by Vegas et al. vegas2015crossover to mitigate these threats at analysis time are effective yet tedious to implement using frequentist methods. However, within Bayesian data analysis, several of the major threats to validity kitchenham2003case can be addressed by including predictors like the period or the interaction effect between the period and treatment in the model. This eases both the mitigation and evaluation of these threats and naturally factors out the impact of confounding factors caused by the design. Bayesian data analysis promises to be an eligible and accessible tool to evaluate data obtained from crossover-design experiments, harnessing the unique advantages of this design for software engineering research vegas2015crossover .
@@ 5.4.1 Internal validity
The crossover design emits additional threats to validity vegas2015crossover . We mitigate the learning by practice effect—i.e., participants getting better when repeating the experimental task—in three ways: Firstly, we disperse the learning effect evenly at design time by randomizing the sequences of treatments. Secondly, we include a warmup object to get participants used to the task and tool but exclude that data from the analysis. Thirdly, we include the period variable as a predictor to factor out the learning effect during the analysis. We avoid the threat of copying by prohibiting communication among participants and using experimental objects where solutions cannot be copied from one task to another.
The threat of optimal sequence describes the risk that there is a sequence in which the treatment is applied, which optimizes the participants’ performance in deriving domain models. We cannot block this threat at analysis time as the sequence and participant IDs are highly correlated. This is because we could—in all but one cases—assign only one participant ($n_{p}=25$) to each sequence ($n_{s}=n_{r}!=4!=24$). Because of this strong correlation, the Bayesian model is incapable of distinguishing between the impact of the sequence ($\beta_{seq}$) from the within-participant variance ($\alpha_{PID}$vegas2015crossover . More participants per sequence would have been necessary to block the threat of an optimal sequence, but these were unavailable to us.
@@ 5.4.2 Conclusion validity
We addressed the reliability of measures threat by peer-reviewing the extraction of the dependent variables from the collected domain models, for which we report an inter-rater agreement of 87.23%. We addressed the random heterogeneity of the subjects by a design in which each participant acts as their own control group. Moreover, we focused on including and analyzing context variables related to the participants’ experience. Our sample of participants is not representative of all context factors. Consequently, our Bayesian data analysis cannot identify all causal effects of some context factors. However, by including them in the causal considerations, the effect of the factors is isolated from the potential confounding variables leven2022broken . The conclusion validity of our study is strengthened by applying two different data analysis approaches and comparing their results. The data analysis suffers from the threat of low statistical power when it comes to evaluating interaction effects, as reliably identifying them requires a larger sample size gelman2018interaction . We limit the number of interaction effects considered in our models and discuss the uncertainty around the coefficient estimates to minimize this threat.
@@ 5.4.3 Construct validity
Our study can suffer from mono-operation bias as we focus only on a subset of quality factors that can potentially exist montgomery2022empirical ; frattini2022live . Nevertheless, our goal with this replication is to extend the initial quality factor of passive voice reported by Femmer et al. femmer2014impact to a second one—ambiguous pronouns—which is widespread as indicated by the literature frattini2022live ; montgomery2022empirical . Similarly, a confounding of constructs and level of constructs could influence the outcomes of our study. For example, the presence of several ambiguous pronouns or passive voice sentences rather than their binary presence or absence from a specification. Further replications, focusing on improving construct validity, should include several levels of each treatment. Mono-method bias is a potential threat to construct validity—i.e., we measured the dependent variables using a single type of measurement, following the original study. However, the measurements were based on a pre-defined protocol and peer-reviewed. Our study may result in a restricted generalizability across constructs since the presence or absence of the different quality factors could result in side effects for other interesting outcomes we did not measure (e.g., comprehensibility or maintainability of the specification). Among the social threats to construct validity, we acknowledge that hypothesis guessing may have taken place since the participants could try to guess the concrete goal of the experimental task based on the invitation text and material provided during the sessions. Nevertheless, we used the same text and phrasing to invite all participants and the same material during the experimental task. Finally, evaluation apprehension could have played a role since some participants are students at the authors’ institution. However, students did not receive rewards (e.g., extra grade points) for participating in the experiment, and they received their course grades before the start of the experiment.
@@ 5.4.4 External validity
The main threat to the external validity of this study is the interaction of setting and treatment as the size and the complexity of the selected specifications, despite being sampled from a real-world data set, might not be representative of the industrial practice.
@@ 6 Conclusion
Requirements quality research lacks empirical evidence and research strategies to advance beyond proposing and following normative rules with unclear impact frattini2022live to better understanding and solving problems relevant to practice franch2017practitioners ; franch2020practitioners . In the scope of our study, we replicated the only controlled experiment on the impact of requirements quality defects on subsequent activities. We demonstrated a method of evaluating data collected through a controlled experiment using a crossover design with Bayesian data analysis. We showed the impact (1) of requirements quality defects varies and (2) may be mediated by context and confounding factors.
@@
We envision that adopting sophisticated statistical tools like Bayesian data analysis and the focus of empirical studies on investigating the impact of requirements quality defects will steer requirements quality research in a relevant and effective direction. We hope that the documentation of this study inspires fellow researchers to adopt our method and tools for replication.
@@ Acknowledgements.
This work was supported by the KKS foundation through the S.E.R.T. Research Profile project at Blekinge Institute of Technology. We are deeply grateful to Parisa Yousefi from Ericsson AB for recruiting practitioners to the experiment.
@@
All supplementary material, including protocols and guidelines for data collection and extraction, the raw data, analysis scripts, figures, and results, are available in our replication package. The replication package is available on GitHub at https://github.com/JulianFrattini/rqi-proto and archived on Zenodo at https://doi.org/10.5281/zenodo.10423666.

### v3

@@ 5.1.1 Answer to RQ1
Using passive voice in natural language requirements specifications has a slightly negative effect on the domain modeling activity regarding missing associations. This finding aligns with the conclusions drawn by the original study by Femmer et al. femmer2014impact ; frattini2024second . However, the Bayesian data analysis emphasizes that both context factors, and especially the number of missing entities, have a significant impact on the number of missing associations as well. Overall, these results support the claim that passive voice can have a negative impact in specific cases but is overall not a significant factor in subsequent activities depending on the requirement femmer2014impact ; krisch2015myth .
@@ 5.2 Implications for Requirements Quality Practice
Finally, the potential influence of context factors on the impact of quality defects on affected activities may incentivize organizations to invest in developing these factors. For example, improving domain knowledge and providing formal modeling training may compensate for quality defects.
@@ 5.3 Implications for Requirements Quality Research
Employing Bayesian data analysis to investigate the impact of requirements quality defects provides sophisticated and sensitive insights necessary to propel requirements quality research frattini2023requirements . The result of the analysis models both the direction and strength of an impacting factor while retaining information about its certainty. These insights go beyond the point-wise comparison and binary result of frequentist analyses furia2019bayesian . The frequentist analysis fails to compare the impact of quality defects on the response variables, as even the calculated effect sizes are similar ($0.930.79<|ES|<$ 0.93). The Bayesian data analysis, on the other hand, clearly shows that some effects are much stronger (e.g., $H_{0}^{AP\rightarrow A^{\times}}$) than others (e.g., $H_{0}^{PV\rightarrow A^{-}}$). Still, the BDA relies on the causal model expressed in a DAG, statistical assumptions about variable types and their independence, and the validity of constructs. Therefore, the results obtained via BDA cannot be seen as more valid by design. However, the BDA is more transparent and allows critical debate—e.g., about the causal assumptions underlying our analysis in Figure 10—which facilitates the incremental improvement of empirical studies.
Including context factors in the prediction allows comparing the impact of requirements quality with the impact of human and process factors, revealing which causes changes in the response variable. These context factors can also represent the properties of non-human agents like generative artificial intelligence (GenAI) models which are increasingly employed for RE tasks. Involving context factors like the version number of a GenAI model, its parameters, its context window, and other factors in empirical studies resembling our approach will allow to investigate which configurations of these models excel at performing their RE task.
Abandoning simple NHSTs for identifying relevant factors of requirements quality and instead opting for a proper framework for causal inference like BDA will increase the likelihood of solving problems with practical relevance franch2020practitioners that justify subsequent tool development frattini2024nlp4re . Empirical studies with explicit causal assumptions (e.g., visualized as DAGs) and sophisticated analyses will produce context-sensitive evidence that can be synthesized in the common framework of the requirements quality theory frattini2023requirements . The continuous synthesis of evidence from individual studies in this common framework will produce more reliable and generalizable conclusions badampudi2019contextualizing ; ciolkowski2005accumulation and effectively address the lack of empirical insights in requirements quality research montgomery2022empirical .
@@
Using a controlled experiment benefits the investigation of the quality factor femmer2014impact . The DAG shown in Figure 10 visualizes this control, as no other factor influences the treatment in question. This eliminates spurious associations that could confuse the results mcelreath2020statistical . On the other hand, the cost of conducting a controlled experiment—especially with participants from industry—cannot be neglected sjoberg2003challenges . Luckily, statistical causal inference via Bayesian data analysis works equally well with observational data, as shown by Furia et al. furia2023towards .
@@
Worth noting is that our comparison between FDA and BDA conflates the use of causal frameworks with advanced Bayesian statistics. An FDA can also employ causal frameworks that mitigate parts of the shortcomings mentioned in Section 2.2, as previously shown by Furia et al. furia2023towards . However, frequentist approaches tend to limit their analyses to the treatment and the response variable, disregarding potential context mund2015does or experimental design factors vegas2015crossover . BDA, on the other hand, entails the use of an explicit causal framework mcelreath2020statistical ; pearl1995bayesian , which is why we support the recommendation of abandoning FDA for BDA in SE research torkar2020missing ; furia2019bayesian .
Implications Quality defects in requirements specifications have a varying impact on affected activities that depend on them. Context factors may compensate for this impact but require better metrics to quantify them. Bayesian data analysis provides more fine-grained insights into these effects than frequentist methods.
@@ 5.4.1 Internal validity
The crossover design emits additional threats to validity vegas2015crossover . We mitigate the learning by practice effect—i.e., participants getting better when repeating the experimental task—in three ways: Firstly, we disperse the learning effect evenly at design time by randomizing the sequences of treatments. Secondly, we include a warm-up object to get participants used to the task and tool but exclude that data from the analysis. Thirdly, we include the period variable as a predictor to factor out the learning effect during the analysis. We avoid the threat of copying by prohibiting communication among participants and using experimental objects where solutions cannot be copied from one task to another.
The threat of optimal sequence describes the risk that there is a sequence in which the treatment is applied, which optimizes the participants’ performance in deriving domain models. We cannot block this threat at analysis time as the sequence and participant IDs are highly correlated. This is because we could—in all but one case—assign only one participant ($n_{p}=25$) to each sequence ($n_{s}=n_{r}!=4!=24$). Because of this strong correlation, the Bayesian model is incapable of distinguishing between the impact of the sequence ($\beta_{seq}$) from the within-participant variance ($\alpha_{PID}$vegas2015crossover . More participants per sequence would have been necessary to block the threat of an optimal sequence, but these were unavailable to us.
@@ 5.4.2 Conclusion validity
We addressed the reliability of measures threat by creating and disclosing evaluation guidelines and peer-reviewing the extraction of the dependent variables from the collected domain models. Despite the acceptable inter-rater agreement score, an in-depth qualitative evaluation of the remaining disagreements may be beneficial to further improve the evaluation instrument and, therefore, the reliability of the results. We addressed the random heterogeneity of the subjects by a design in which each participant acts as their own control group.
Moreover, we focused on including and analyzing context variables related to the participants’ experience. Our sample of participants is not representative of all context factors. Consequently, our Bayesian data analysis cannot identify all causal effects of some context factors. However, by including them in the causal considerations, the effect of the factors is isolated from the potential confounding variables leven2022broken .
The conclusion validity of our study is strengthened by applying two different data analysis approaches and comparing their results. The data analysis suffers from the threat of low statistical power when it comes to evaluating interaction effects, as reliably identifying them requires a larger sample size gelman2018interaction . We limit the number of interaction effects considered in our models and discuss the uncertainty around the coefficient estimates to minimize this threat.
The analysis can suffer from violated assumptions of statistical tests. Modeling the number of missing entities and associations as binomial distributions implies the independence of each event, i.e., that each missing entity and association is independent of all other missing entities and associations. While we did not observe any cascading, i.e., dependent, defects, their independence remains only assumed.
@@ 5.4.3 Construct validity
Our study can suffer from mono-operation bias as we focus only on a subset of quality factors that can potentially exist montgomery2022empirical ; frattini2022live . Nevertheless, our goal with this replication is to extend the initial quality factor of passive voice reported by Femmer et al. femmer2014impact to a second one—ambiguous pronouns—which is widespread as indicated by the literature frattini2022live ; montgomery2022empirical .
Similarly, a confounding of constructs and level of constructs could influence the outcomes of our study, for example, the presence of several ambiguous pronouns or passive voice sentences rather than their binary presence or absence from a specification. Further replications, focusing on improving construct validity, should include several levels of each treatment.
Mono-method bias is a potential threat to construct validity—i.e., we measured the dependent variables using a single type of measurement, inspired by the original study. However, the measurements were based on a pre-defined protocol and peer-reviewed. Our study may result in a restricted generalizability across constructs since the presence or absence of the different quality factors could result in side effects for other interesting outcomes we did not measure (e.g., comprehensibility or maintainability of the specification).
Among the social threats to construct validity, we acknowledge that hypothesis guessing may have taken place since the participants could try to guess the concrete goal of the experimental task based on the invitation text and material provided during the sessions. Nevertheless, we used the same text and phrasing to invite all participants and the same material during the experimental task. Evaluation apprehension could have played a role since some participants are students at the authors’ institution. However, students did not receive rewards (e.g., extra grade points) for participating in the experiment, and they received their course grades before the start of the experiment.
Moreover, our study can suffer from an inadequate preoperational explication of constructs as we did not validate our context factors. For example, we are unable to provide any proof that the self-reported number of years spent in RE adequately represents the latent variable of experience in RE beyond educated guesses and relying on comparable practices in our scientific community bogner2023restful . To improve the construct validity, separate studies investigating the adequacy of these measurements in representing their constructs are necessary. This particularly impacts our decision to replace the binary distinction of participants by type (students versus practitioners) with more fine-grained variables like experience and domain knowledge. While our study supports the feasibility of this step on an analytical level, we cannot prove its validity on a conceptual level. We encourage investigating the feasibility of variables to represent individual skills to improve the construct validity of studies considering this impact salman2015students .
Finally, a variable of the selected population that may interact with the treatment that we did not analyze is the language skill of participants. Arguably, skills in the English language influence the ability to comprehend and process the experimental objects and, therefore, may impact the response variables. We were unable to measure this variable properly given that all participants scored the same on the Common European Framework of Reference for Languages (CEFR) martyniuk2006common (i.e., non-native, fluent English speakers). While the threat is minimized in our study due to the comparable language level of participants, future studies should develop measurement instruments for this construct and involve this variable in such causal queries.
@@ 5.4.4 External validity
The main threat to the external validity of this study is the interaction of setting and treatment as the size and the complexity of the selected specifications, despite being sampled from a real-world data set, might not be representative of the industrial practice. Using Google Docs as the modeling tool is not fully representative of real-world practices. Given that it was appropriate and sufficient for the experimental task, however, renders this as an opportunity for improving the realism of the experiment in future studies rather than a threat to validity.
There may be the threat of interaction of selection and treatment, as some participants reported no modeling experience or training. These deficiencies might influence the results and render a subset of the participants as non-representative of our target population. We attempted to mitigate this threat via comprehensive instructions and including a warm-up phase in the experiment.
@@ 6 Conclusion
Requirements quality research lacks empirical evidence and research strategies to advance beyond proposing and following normative rules with unclear impact frattini2022live to better understanding and solving problems relevant to practice franch2017practitioners ; franch2020practitioners . In the scope of our study, we conducted a controlled experiment on the impact of requirements quality defects on subsequent activities. We demonstrated a method of evaluating data collected through a controlled experiment using a crossover design with Bayesian data analysis. We showed the impact (1) of requirements quality defects varies and (2) may be mediated by context and confounding factors. The part of our study that serves as a conceptual replication strengthens the claims of the re-analyzed original study femmer2014impact ; frattini2024second that passive voice only has a slight impact on missing associations from domain models.
@@
We envision that adopting sophisticated statistical tools like Bayesian data analysis and the focus of empirical studies on investigating the impact of requirements quality defects will steer requirements quality research in a relevant and effective direction. Explicit causal assumptions and sophisticated data analyses will produce empirical evidence which can be more easily synthesized to more reliable and generalizable conclusions ciolkowski2005accumulation in a common framework frattini2023requirements . We hope that the documentation of this study inspires fellow researchers to adopt our method and tools for replication.
@@ Acknowledgements.
This work was supported by the KKS foundation through the S.E.R.T. Research Profile project at Blekinge Institute of Technology. We are deeply grateful to Parisa Yousefi from Ericsson AB for recruiting practitioners to the experiment. We further thank the reviewers for their tremendous effort that significantly improved this manuscript.
@@
All supplementary material, including protocols and guidelines for data collection and extraction, the raw data, analysis scripts, figures, and results, are available in our replication package frattini2024replication . The replication package is available on GitHub at https://github.com/JulianFrattini/rqi-proto and archived on Zenodo at https://doi.org/10.5281/zenodo.10423666.